### v1

# Right Reward Right Time for Federated Learning

Abstract
Critical learning periods (CLPs) in federated learning (FL) refer to early stages during which low-quality contributions (e.g., sparse training data availability) can permanently impair the learning performance of the global model owned by the model owner (i.e., the cloud server). However, strategies to motivate clients with high-quality contributions to join the FL model training process and share trained model updates during CLPs remain under-explored. Additionally, existing incentive mechanisms in FL treat all training periods equally, which consequently fails to motivate clients to participate early. Compounding this challenge is the cloud’s limited knowledge of client training capabilities due to privacy regulations, leading to information asymmetry. Thus, in this article, we propose a time-aware incentive mechanism, called Right Reward Right Time (R3T), to encourage client involvement, especially during CLPs, to maximize the utility of the cloud server in FL. Specifically, the cloud utility function captures the trade-off between the achieved model performance and payments allocated for clients’ contributions, while accounting for the client’s time and system capabilities, the efforts, joining time, and rewards. Then, we analytically derive the optimal contract for the cloud and devise a CLPs-aware mechanism to incentivize early participation and efforts while maximizing cloud utility, even under information asymmetry. By providing the right reward at the right time, our approach can attract the highest-quality contributions during CLPs. Simulation and proof-of-concept studies show that R3T increases cloud utility and is more economically effective than benchmarks. Notably, our proof-of-concept results demonstrate up to a 47.6% reduction in the total number of clients and up to a 300% improvement in convergence time while reaching competitive test accuracies compared with incentive mechanism benchmarks.
Index Terms:
Blockchain, Contract theory, Critical learning periods, Federated learning, Incentive mechanism.
1 Introduction
1.1 Background and Motivations
Federated learning (FL) is a privacy-preserving data sharing and learning paradigm, where distributed clients (e.g., devices or organizations) collaboratively train an artificial intelligence model without disclosing their private data [1, 2]. In FL settings, the importance of all training rounds and the efforts from distributed clients across these rounds are regarded as equally important. However, this setting is questioned by a recent discovery of a phenomenon in FL known as critical learning periods (CLPs), which highlight the varying impacts of different training rounds and significantly affect FL learning and training efficiency [3]. This finding was inspired by earlier research on CLPs in species’ cognitive and learning functions (i.e., humans and animals) and deep learning networks within a centralized setting [4, 5, 6, 7]. Particularly, CLPs are early learning phases (c.f., Fig. 1(1(a))) profoundly affect long-term learning capabilities [8]. Empirical studies show that the absence of essential and proper learning experiences during these periods can lead to irreversible deficits. For instance, barn owls exposed to misaligned auditory and visual cues during their critical development periods fail to properly localize spatial locations in adulthood [9]. This phenomenon was also observed in deep neural network, in which any deficits and low effort (e.g., sparse data availability, late client participation) during this initial learning stage can cause a lasting impaired effect on learning model performance, regardless of subsequent efforts (c.f., Fig. 1(1(b))). The reason for the phenomenon is still unclear, but possible explanations include the depth of the learning model, structure of training data distribution, or defects in the implementation and training process [5, 6, 7].
Figure 1: Illustration of: a) CLPs in the biological field [10], and b) how early effort affects long-term FL model performance.
Despite the promising merits of enhancing FL learning and training efficiency (e.g., accuracy and convergence rate), CLPs in FL remain under-explored, including a lack of transparent, fair, and timely incentives for participation and contribution enticement and information asymmetry among clients and the global model owner (e.g., cloud server) in these periods. Current existing CLP-considered works in FL, such as CLPs detection, client selection, or CLP-aware robustness, are based on a common assumption of voluntary engagement of clients in the training process (e.g., [3, 11, 12, 13]). However, this is impractical because clients are rational, strategic, and self-interested and need sufficient rewards to compensate for their efforts (e.g., data, computation resources, and time costs). This leads to a lack of client motivation for contribution, resulting in the unsustainable development of FL systems. Although several works have been proposed by authors (e.g., [14, 15, 16, 17]) to bring incentivization by considering effort-reward approaches to fairly compensate FL clients’ effort, they treat clients’ efforts and their impacts on the global model performance on every training round equally and ignore the temporal factor, whereas efforts in CLPs have more weights and mainly contribute to the overall performance of the FL global model. By failing to prioritize CLP-driven incentive solutions, current methods inefficiently allocate cloud budgets (e.g., monetary rewards), thereby reducing cloud utility—defined as the difference between global model performance and compensation costs (which is defined later in Section 2).
Additionally, there are significant issues of information asymmetry between FL clients and the cloud server, which make CLPs misaligned and cause disputes among the cloud server [^1] The cloud server and the cloud are used interchangeably. and clients. Specifically, due to privacy restrictions, the cloud server lacks prior knowledge of the clients’ training capabilities, including the computational resources, data properties (e.g., data quantity or distribution), or available joining time. Therefore, assumptions that the cloud server knows precise information [18, 19, 20], such as clients’ computing resources or clients’ data size, may not hold in practice. This uncertainty can adversely affect the alignment of CLPs in determining the final model performance in FL. For example, under conditions of information asymmetry, the cloud server randomly selects clients with different training capacities, resulting in the underrepresentation of high-quality clients [^2] We refer high-quality clients to clients with high-quality contributions and efforts, such as a huge amount of training data and large computation resources, and they are available in CLPs., thereby high-quality contributions and efforts may get diluted with low-quality ones when averaged together, impairing the early advantage impact of CLPs on the final model performance. There are also uncertainty regarding whether clients will exert the promised effort post-contract [21]. Despite the presence of contractually agreed-upon incentives, FL clients may still engage in dishonest behaviour, such as deliberately denying agreed-upon compensation, thus leading to disputes among clients and the cloud server [22]. These incur high costs to the cloud server to monitor such information and may make misjudgments, increasing the difficulty of achieving a sustainable FL system and fair incentive mechanism [15]. In this article, we aim to address the following research question (RQ) in FL.
RQ: "How can we design an incentive mechanism to attract rational clients with highest-quality contributions to improve the FL learning and training efficiency, and mitigate information asymmetry between clients and the cloud server, given the crucial role of CLPs in determining the final FL model performance?"
To address these aforementioned challenges, we propose a CLP-aware incentive mechanism, named Right Reward Right Time (R3T), which leverages contract theory to effectively handle the information asymmetry issues [21]. In addition, R3T integrates with blockchain smart contract to handle the issues of centralized incentive governance and disputes caused by relying on the cloud server [23, 24, 25]. Our approach is supported with theoretical modeling and analysis, together with intensive experiments. Specifically, a set of tailored contract items is designed by the cloud server for clients with different training capabilities per training round. Realizing the importance of CLPs, the cloud server optimally prioritizes and provides corresponding generous rewards for clients who joined early and contributed efforts in the early training phase under complete and incomplete information cases. By providing the right reward at the right time, R3T enables highest-quality clients to join early and contribute efforts for learning-based tasks given by the existence of CLPs in FL thanks to open, transparent, and fair contract items executed on blockchain smart contracts.
1.2 Contributions
The main contributions of this paper are as follows:
[…]
    CLP-aware incentive mechanism design: We propose a time-aware contract theory-based model, called R3T, to characterize the cloud utility as a function of rational client’s time and system capabilities, effort, joining time, and reward. Then, we develop a CLP-aware incentive solution and algorithm to jointly optimize client strategies and cloud utility while adaptively attracting high-quality rational clients in the early training stages to enhance FL training and learning efficiency. The proposed incentive mechanism provides the right reward at the right time and handles issues of information asymmetries between clients and the cloud server.
[…]
    Performance evaluation: We carry out simulations to verify the effectiveness of our R3T in information symmetry and information asymmetry cases under individual rationality, incentive compatibility, and budget feasibility constraints. R3T demonstrates feasibility and superior efficiency compared with conventional incentive mechanism benchmarks, including linear pricing and contract theory-based methods.
[…]
    Proof of concept: We develop a proof of concept of R3T and demonstrate the feasibility and efficiency of the proposed mechanism and system design. Significantly, we show that R3T can boost the convergence time by up to 300% while achieving competitive final FL model accuracies on standardized benchmark datasets, CIFAR-10 and Fashion-MNIST, compared with the state-of-the-art incentive mechanism benchmarks.
1.3 Related Work
In the following, we discuss related work concerning CLPs in artificial neural networks and incentive mechanisms.
Critical learning periods in artificial neural networks. Achille et al. in [5] empirically investigated the correlation between critical periods in the memorization phase and the training outcome in deep neural networks. Their findings demonstrated that analogous to animals, sensory deficits in CLPs cause lasting impairments to the information and connectivity of the network, as quantified by an approximation of the Fisher Information, and adversely affect learning outcomes regardless of subsequent additional training. Similarly, Kleinman et al. in [26] analytically found that the depth of the deep neural network model and data distribution structure are fundamental sources of CLPs in deep neural networks, not only being explained by ascribing to biological ageing processes observed in animals. A similar pattern was found in the distributed training manner, where Yan et al. in [3, 11] showed that FL exhibits CLPs, which play a vital role in determining the final training performance, given the complex of statistical and system heterogeneity. To further explain the phenomenon, metrics such as federated Fisher Information matrix and CLP-aware federated norm vector are proposed to trace and detect CLPs. However, in FL settings, these CLP-given studies are based on the assumption that clients act without cost compensation, rationality, self-interest, and strategic behavior, which may not accurately reflect real-world conditions. Without proper CLP-aware incentives, it may lead to the lack of high-quality contributions in CLPs that have been shown to be permanent and irrecoverable impairment on the global model performance, despite subsequent efforts.
Incentive mechanism design. By incorporating incentive mechanisms into FL systems, engagement and cooperation are promoted where clients are rational, resulting in enhanced learning and training efficiency [27]. Realizing the importance of incentive mechanisms, there are several efforts to develop game-theoretic incentive techniques in FL (e.g., [28, 29, 30, 31, 32, 33, 34]). Authors in [28, 29, 30] designed leader-follower-based incentive mechanisms to encourage clients’ contribution, where the cloud server sets the total reward or price per contribution unit, and then FL clients determine contribution strategies (e.g., data quantity or local performance) to maximize their utilities. However, the information asymmetry condition was not considered in these studies. To handle this issue, utilizing the self-revelation principle of contract theory [21], Sun et al. in [31] proposed an incentive mechanism using a single client’s property (i.e., privacy leakage costs) to create different contract items under both complete and incomplete information cases to compensate for client’s privacy leakage while ensuring comparable learning model performance. Under the same information cases, Huang et al. in [33] proposed a data quality-based incentive mechanism, where the client’s data size, data quality, and unit data computing cost are considered, to derive optimal contract items to optimize the cloud server’s costs in the era of artificial intelligence-generated. Tang et al. in [34] proposed a non-cooperative game-based incentive mechanism to model the interactions between clients (i.e., organizations), deriving optimal processing capacity of each client, to address free-rider problems of public goods in FL under constraints of individual rationality and budget balance. To provide further comprehensive incentive solutions by handling the transparency, centralization, and potential disputes among clients and the cloud server, blockchain is fused with FL systems to manage incentive calculation and distribution [29, 35, 30, 36, 37, 34]. For instance, Yuan et al. in [37] utilized Ethereum-based smart contracts to autonomously execute an incentive mechanism, that considers influences from competition and provides fair and non-repudiative compensation and incentive, to improve social welfare. Nevertheless, existing works on incentive mechanisms in FL to provide the right reward at the right time have not been explored yet. None of these works considers the existence of CLPs, which have shown to play a vital role in determining final FL model performance, and these studies treat every training round as equally important. Besides, information asymmetries among clients and the cloud server, and their impacts on CLPs’s advantages and the final model performance have not been analyzed yet.
To the best of our knowledge, R3T is the first CLP-aware incentive mechanism, considering the joint importance of joining time and effort, to attract the highest-quality rational clients to participate in the FL model training in early stages and mitigate information asymmetry between clients and the cloud server. R3T provides a transparent and autonomous execution of reward calculation and distribution.
2 System model
2.1 System Overview
Figure 2: R3T system’s learning workflow: (1) prepare contracts, (2) sign contracts & download global model, (3) local training, (4) upload updates, (5) forward & validate updates, (6) receive & aggregate model updates, and (7) settlement.
As depicted in Fig. 2, the R3T system consists of three main subsystems: edge servers forming a blockchain platform, clients, and a cloud server. We denote a set of edge servers as $\mathcal{L}$, where each edge server $l\in\mathcal{L}$ connects with clients to exchange information. $\mathcal{L}$ forms, maintains, and operates a blockchain platform to record and validate information such as client identification (ID), local model updates, global models, joining time, and effort levels. Also, R3T integrates a distributed file storage system such as InterPlanetary File System (IPFS), to achieve privacy preservation, high availability, scalability, and low latency for data storage and retrieval [38]. The set of clients is denoted as $\mathcal{N}=\{1,...,n,...,N\}$, where $N$ represents the number of clients. Clients join and contribute their time, computing, and data resources to jointly train the global model published by the cloud server, in exchange for benefits such as monetary rewards [39]. In each training round, a subset of clients $\mathcal{S}_{n}\subseteq\mathcal{N}$, participates in improving the model performance (e.g., in the training round $t$, the cloud server randomly selects $\mathcal{S}_{n}^{t}$ clients out of $\mathcal{N}$). In addition, each client owns or has a right to use or share its dataset $\mathcal{D}$, and the local loss function of a client is defined as ${F}$. In the system, the cloud server plays roles of initiating the global model parameters ${w}$ and seeking high-quality clients to optimize the model performance. It receives local model updates from clients, conducts model aggregation, and handles payments via a blockchain platform.
2.2 Learning Workflow
R3T is designed to perform model training, storage, and reward calculation and allocation in a transparent, fair, and auditable manner, leveraging blockchain and FL technologies. It integrates a time-aware contract theory-based mechanism to address right-time and right-reward allocation where CLPs phenomenon exists. The learning process of R3T involves the following steps. 1) Prepare contracts. The cloud server initiates its naive FL model ${w}$, designs a set of contract items, and distributes them to $N$ registered clients. 2) Sign contracts & Download global model. Each client signs a contract item to participate in the training process and downloads the global model ${w}$. 3) Local training. Clients train ${w}$ using their private data $\mathcal{D}$. 4) Upload updates. Clients send their trained model updates, metadata (e.g., IDs, addresses, chosen contracts), and hashed trained model updates to their associated edge servers after training [^3] We assume that clients are truthful in reporting their updates to the cloud. If required, the cloud can verify their updates using the Trusted Execution Environments [40] or reputation systems [41]. Besides, given the traceability property of the underlying blockchain [24] and incentive mechanism design (e.g., incentive compatibility) presented later in Section 3, clients are encouraged to truthfully report to benefit from joining and contributing.. 5) Forward & Validate updates. Edge servers relay trained model updates to the cloud server to prevent delay in global model aggregation process and model convergence rate, and only cross-validate hashed trained model updates and metadata to ensure data integrity and maintain consensus, storing relevant data on the blockchain for transparency and arbitration in case of disputes. 6) Receive & Aggregate model updates [^4] To avoid the straggler effect, we employ a synchronous learning update scheme, which has a provable convergence [42, 27].. The cloud server aggregates them using an aggregation algorithm such as FevAvg [1]. 7) Settlement. The process iterates until the global model converges. Upon completion, the final global model is transmitted back to the cloud server, which issues rewards based on the signed contract items. The rewards, calculated using blockchain smart contracts, are distributed to clients to compensate for their contributions (i.e., time and effort) to the learning task. Edge servers are also incentivized by receiving rewards for validating and storing transaction data.
2.3 Time Frame, Client Type, Contract, and Strategy
1) Time Frame: The time frame is divided into $T$ training rounds [^5] In this work, training rounds and time slots are used interchangeably., denoted as $\mathcal{T}=\{1,...,t,...,T\}$, in which CLPs are vital duration in determining the final model performance.
2) Client Type: $N$ clients are classified into a set $\mathcal{K}=\{1,...,k,...,K\}$ of $K$ distinct types according to their available time and system capabilities (i.e., computation and communication resources, and the quantity of data) due to data and system heterogeneity among them in real-world conditions. We denote the capability of a client $n$ in terms of time and system resources of a type-$k$ client ($1\leq k\leq K$) by ${\theta}_{n,k}$ or ${\theta}_{k}$ [^6] For simplicity, we interchangeably represent ${\theta}_{n,k}$ and ${\theta}_{k}$ for a type-$k$ client $n$’s capability. Similarly, the contract item $\phi_{k}$ can be denoted as $\{{e_{n,k},t_{n,k},R_{n,k}}\}$ or $\{{e_{k},t_{k},R_{k}}\}$. The same notation applies to the model parameter $w$, which is introduced later..
Definition 1.
Without loss of generality, client training capabilities in $K$ types belong to a discrete and finite set, as sorted in ascending order by $0<{\theta}_{1}<...<{\theta}_{K}.$
It signifies that higher client types have greater training capabilities and a higher willingness to participate in early-available time, especially in CLPs, as well as provide a huge amount of data. The cloud is unaware of the exact client types, knowing only the probability of a client belonging to a certain type $k$. Due to privacy regulations, direct collection of client-specific information is restricted, making it challenging to predict individual client capabilities. Instead, only aggregate statistical insights can be obtained through market research and surveys conducted with client consent [43].
3) Cloud’s Contract: To motivate clients to join early, share resources, and conduct training, the cloud must offer transparent and fair contribution-reward bundles, compensating clients for their efforts. Recognizing the importance of CLPs, the cloud offers extra bounties to attract high-quality, resource-rich clients to participate actively. Consequently, each client’s gain is influenced by their chosen time of participation in the training process. The cloud has a contract set $\bm{\phi}=\{\phi_{k}\}_{k\in\mathcal{K}}$, which contains $K$ contract items corresponding to $K$ client types. The contract item $\phi_{k}=\{e_{k},t_{k},R_{k}\}$ specifies the relationship between type-$k$ client’s effort, joining time, and reward. Specifically, $e_{k}$ is defined as the number of data units required for each training round, and $t_{k}$ is the training time slot, which is either in CLPs or in non-CLPs. Also, $R_{k}$ is the corresponding reward for type-$k$ client paid by the cloud server for finishing the training round, and is given as follows
$R_{k}=r_{k}+B_{k},B_{k}=\theta_{k}h(t_{k})e_{k},$ (1)
where $r_{k}$ is the basic salary and $B_{k}$ is the bonus, which has been shown as an effective payment structure to motivate clients’ efforts toward collective goals (i.e., enhancing global model performance) [44]. Specifically, clients are assured to have a base compensation of salary $r$ upon joining the training process. Based on their individual types, the timing of their participation, and the level of effort they contribute, bonuses are awarded. These bonuses direct clients participating in CLPs to contribute more efforts during this period, as well as providing additional rewards for the highest-contributing clients [45]. Besides, inspired by the demand-price relationships in [46], we propose a function $h(t_{k})$ which is a time-aware bonus unit function (i.e., bonus-time relationship function), where the crucial role of CLPs is known by clients, and is presented as follows
$h(t_{k})=\left\{\begin{array}[]{ll}1+\frac{\vartheta}{\ln(2t_{k})}&\mbox{if ${t_{k}}\in\text{CLPs}$};\\ 1&\mbox{if ${t_{k}}\in\text{non-CLPs}$},\end{array}\right.$ (4)
where $\vartheta$ represents the bonus unit coefficient. That function offers high-quality clients bonuses when they choose to join the model training process at an early stage, given the importance of CLPs.
4) Client’s Strategy Space: Each client decides whether to participate in the training process, (if yes) how much effort they should exert and which time slot to join. The choice of different time slots may result in varying bonuses. Particularly, participating during the CLPs yields higher payoffs for clients who contribute significant effort, as training the global model with huge effort during these periods is crucial for improving final model performance.
2.4 Utilities
According to (1), $R_{k}$ is dependent of $e_{k}$, $t_{k}$, and $r_{k}$, thus each contract item is rewritten by $\phi_{k}=\{e_{k},t_{k},r_{k}\}$.
1) Clients: The utility of type-$k$ client $n$ can be defined by
$\mathcal{U}_{n}(e_{k},t_{k},r_{k})$ $=R_{k}-\left(\frac{1}{2}\delta e^{2}_{k}+\beta r_{k}\right)$
$={\theta}_{k}h(t_{k})e_{k}-\frac{1}{2}\delta e^{2}_{k}-(\beta-1)r_{k},$ (5)
where $\left(\frac{1}{2}\delta e_{k}^{2}+\beta r_{k}\right)$ is the total contribution cost. This includes $\frac{1}{2}\delta e^{2}_{k}$, which is a quadratic cost function with respect to client’s effort (e.g., sensing, collecting, and training of data or using cloud services for these tasks) according to [21], with $\delta>0$ being a quadratic unit effort cost that varies on different learning tasks (e.g., generating data samples or training data samples). For example, training ChatGPT-3 with 45TB of compressed plaintext before filtering and 570GB after filtering eighty using 80 V100 GPUs costed more than \$2 million or generating 100,000 data points using Amazon’s Mechanical Turk service will cost around \$70,000 [47]. Additionally, $\beta r_{k}$ represents the time cost associated with joining time $t_{k}$, considered as a linear function of the reward, where $\beta>1$ is the time cost coefficient [21, 48].
2) Edge Server: The edge server $l$’s utility is defined as $\mathcal{U}_{l}={R_{l}}-v(f_{l}),$ where $R_{l}$ is the gain for data mining paid by cloud, $v(f_{l})$ is the mining cost function, and $f_{l}$ is the edge server’s computing resource such as CPU cycle frequency.
3) Cloud server: The cloud’s utility is the difference between the gain (i.e., FL model performance) and costs (i.e., rewards for clients’ contributions such as joining time and efforts in training the FL model, and for edge server in mining and validating data transactions) as follows:
$\mathcal{U}_{c}(\bm{e},\bm{t},\bm{r})=\lambda\sum_{k\in\mathcal{K}}g(h(t_{k})e_{k})-\sum_{k\in\mathcal{K}}{R_{k}}-\sum_{l\in\mathcal{L}}{R_{l}},$ (6)
where $\lambda$ is a parameter adjusted by the cloud server to show its concerns for FL model performance (i.e., larger $\lambda$) or reward budget (i.e., smaller $\lambda$), $\bm{e}=\{e_{1},\dots,e_{K}\}$, $\bm{t}=\{t_{1},\dots,t_{K}\}$, and $\bm{r}=\{r_{1},\dots,r_{K}\}$. Without the loss of generality, $g(\cdot)$ is defined as a concave function [14, 49] with respect to the amount of effort and joining time. Specifically, with the same effort, each client can make different contributions and impacts on FL model performance at different training times.
2.5 Blockchain smart contract-based R3T procedure
Our focus is on the incentive mechanism design given the existence of CLPs in FL, which can be considered as a service and is built on top of the underlying blockchain [23, 24, 50], thereby issues of blockchain security and scalability are out of scope in this article. Specifically, R3T uses smart contracts to credibly enable client and cloud registration, effort record, data storage, and reward calculation and distribution without involvement of third parties. In case of disputes between clients and the cloud, the immutably recorded results on the blockchain can serve as reliable evidence for arbitration.
Figure 3: Procedure of R3T based smart contracts.
Fig. 3 illustrates the functionality of blockchain and smart contracts in the R3T, including four primary operations. 1) First, clients and the cloud server initiate their participation. During this process, the system records relevant information, such as IDs, and wallet addresses, for both clients and the cloud server to ensure that only registered participants can interact within the system. Additionally, the cloud server publishes a contract set $\phi^{*}$, where $\phi^{*}$ is a set of optimal contract items determined by the cloud server in Section 3, directly onto the blockchain for transparency and accessibility to clients. 2) Upon completing the training process, clients upload their strategy profiles or selected contract items $\phi_{k}^{*}$ with their IDs, the location of the model weights, and hashed model weights to the blockchain by invoking contributitonSubmit() function in the Contribution smart contract. 3) To enhance the scalability, minimize on-chain storage overhead, and provide robustness to single-point-of-failure attacks for the cloud server of R3T, dataStore() in the Contribution smart contract is employed. Global model weights are uploaded to private IPFS built by authorized edge servers while the resulting hash and corresponding training round number are recorded on-chain, ensuring data integrity and efficient and decentralized model management. 4) Finally, contract-based rewards $R_{k}$ (e.g., tokens or cryptocurrencies such as Ether in the Ethereum blockchain platform) are calculated and distributed through reward() function in the Reward smart contract.
In the next section, we analyze R3T under two information conditions over CLPs and non-CLPs. 1) Complete Information Case: Clients’ types are public and known to the cloud server and other clients. However, this case is impractical due to privacy leakage risks from exposing individual properties. 2) Incomplete Information Case: Clients’ types are private and only measured locally, but the cloud server may statistically know the distribution of clients’ types through historical market research or surveys [43].
3 R3T - The Proposed Incentive Mechanism for critical learning periods
3.1 Problem Formulation
A feasible contract set ensures clients, who choose their true preference types, receive equitable rewards that are commensurate with their costs and effort. Besides, in each training round, the cloud server can only afford a given budget $P$ for rewarding. To achieve this, it must satisfy the following individual rationality (IR), incentive compatibility (IC), and budget feasibility (BF) constraints:
Definition 2 (Individual Rationality).
Each type-$k$ client only chooses the contract item if and only if its utility is non-negative,
$\mathcal{U}_{n}(e_{k},t_{k},r_{k})\geq 0,\forall{k\in\mathcal{K}},\forall{t_{k}\in\mathcal{T}}.$ (7)
Definition 3 (Incentive Compatibility).
Each type-$k$ client maximizes its utility by choosing the contract item $\{e_{k},t_{k},r_{k}\}$, which is specifically designed for its type and is presented by
$\mathcal{U}_{n}(e_{k},t_{k},r_{k})\geq\mathcal{U}_{n}(e_{k^{\prime}},t_{k^{\prime}},r_{k^{\prime}}),\forall{k,k^{\prime}\in\mathcal{K}},k\neq k^{\prime}.$ (8)
Assumption 1 (Rationality).
Each client is willing to join and contribute its effort to the training process, which guarantees IR and IC constraints and leads to the best utility for itself [21]. For example, realizing the importance of CLPs with a correspondingly huge amount of rewards announced by the cloud server [^7] That is considered as common knowledge [21]., each client tends to join and allocate efforts in this period.
Definition 4 (Budget Feasibility).
The total rewards for all participating type-$k$ clients do not exceed cloud budget $P$, i.e.,
$\sum\nolimits_{k=1}^{K}R_{k}\leq P.$ (9)
To maximize the utility, the cloud server offers the optimal contract set $\bm{\phi^{*}}=\{\bm{e^{*}},\bm{t^{*}},\bm{r^{*}}\}$ to clients. The optimal contract set is the solution to the following problem [^8] The total blockchain mining rewards are constant because the computing resources required for this task are the same across all edge servers in each global training round [35]. Thus, this factor does not affect the optimization and analysis in this paper, so we omit it here.:
$\max_{(\bm{e},\bm{t},\bm{r})}\quad$ $\mathcal{U}_{c}{(\bm{e},\bm{t},\bm{r})},$ (10a)
[…]
Given the cloud server’s contract, we model the game among clients as follows
  • Players: $|\mathcal{S}_{n}|$ clients are in the set $\mathcal{N}$ per training round.
  • Strategies: each client $n\in\mathcal{S}_{n}$ decides which contract item $\phi_{k}=\{e_{k},t_{k},r_{k}\}\in\phi$ to choose based on its capability for model training task published by the cloud server.
  • Objectives: each type-$k$ client $n$ aims to maximize its utility $\mathcal{U}_{n}(e_{k},t_{k},r_{k})$ expressed in (2.4).
Once $t_{k}$ and $r_{k}$ are given, we determine that the utility of a client in (2.4) is a strictly concave function with respect to its effort $e_{k}$. Hence, the optimal choice of effort $e_{k}$ for a type-$k$ client can be obtained by setting the first-order derivative of the client’s utility function with respect to $e_{k}$. Specifically,
$e^{*}_{k}=\frac{\theta_{k}h(t_{k})}{\delta}.$ (11)
From (4) and (11), we can see that the optimal effort of each type-$k$ client is greater than zero and independent of $r_{k}$, but is increasing with its time and system capability $\theta_{k}$ and is decreasing with joining time $t_{k}$. In other words, if a client is willing to join early and has high-system capabilities, it has a higher chance of receiving more bonuses and rewards.
Substituting $e^{*}_{k}$ into (1), (2.4), (6), the utilities of a type-$k$ client and the cloud server are rewritten respectively as
$R_{k}=r_{k}+\frac{\theta_{k}^{2}h^{2}(t_{k})}{\delta},$ (12)
$\mathcal{U}_{n}(t_{k},r_{k})=\frac{{\theta}_{k}^{2}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k},$ (13)
$\mathcal{U}_{c}(\bm{t},\bm{r})=\sum_{k=1}^{K}{\left[\lambda g\left(\frac{\theta_{k}h^{2}(t_{k})}{\delta}\right)-\left({r_{k}+\frac{\theta_{k}^{2}h^{2}(t_{k})}{\delta}}\right)\right]}.$ (14)
The optimal contract design can be rewritten by
$\max_{(\bm{t},\bm{r})}\quad$ $\mathcal{U}_{c}{(\bm{t},\bm{r})},$ (15a)
[…]
3.2 R3T Incentivization under Complete Information Case
We study scenarios where the cloud server knows clients’ time and system capabilities precisely before any interaction. This serves as an ideal benchmark, but it may be infeasible in real-world applications due to privacy regulations.
With precise knowledge of each client’s type (i.e., client’s available time and system capabilities), the cloud server can tailor personalized contracts to clients. Each type-$k$ client receives a corresponding contract item $\phi_{k}=\{e_{k},t_{k},r_{k}\}$, ensuring that IC and BF constraints in (8) and (9) are satisfied. The cloud server only considers the IR constraints in (7) for feasible contract design. The optimal contract design under the complete information case can be formulated as
$\max_{(\bm{t},\bm{r})}\quad$ $\sum_{k=1}^{K}{\left[\lambda g\left(\frac{\theta_{k}h^{2}(t_{k})}{\delta}\right)-{r_{k}-\frac{\theta_{k}^{2}h^{2}(t_{k})}{\delta}}\right]},$ (16a)
s.t. $\frac{{\theta}_{k}^{2}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}\geq 0,$ (16b)
$\sum_{k=1}^{K}\left[r_{k}+\frac{\theta_{k}^{2}h^{2}(t_{k})}{\delta}\right]\leq P,{\forall t_{k}\in\mathcal{T}},{\forall k\in\mathcal{K}}.$ (16c)
Theorem 1.
All optimal contract items satisfy the condition $\frac{{\theta}_{k}^{2}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}=0,\forall{k\in\mathcal{K}}$ under the complete information case, that is the client’s utility is zero.
Proof.
A rational cloud server aims to exploit contributions (i.e., effort and early-available time) from clients, leading to zero utility for clients. To be more specific, the cloud server always chooses a larger effort $e_{k}$ (i.e., a larger $h(t_{k})$ where $e_{k}=\frac{\theta_{k}h(t_{k})}{\delta}$ shown in (11)), and a larger salary $r_{k}$. By increasing these variables, the cloud incentivizes clients to join promptly and invest time and effort in model training. In other words, the cloud exhausts its budget to quickly achieve model convergence as long as the IR constraints are not violated, specifically, $\frac{{\theta}_{k}^{2}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}=0$. ∎
It is important to note that a larger $h(t_{k})$ means a smaller $t_{k}$ that is at the early training stage. The optimal contract design in (16a) is rewritten by
$\max_{(\bm{t},\bm{r})}\quad$ $\sum_{k=1}^{K}{\left[\lambda g\left(\frac{\theta_{k}h^{2}(t_{k})}{\delta}\right)-r_{k}-\frac{\theta_{k}^{2}h^{2}(t_{k})}{\delta}\right]},$ (17a)
s.t. $\frac{{\theta}_{k}^{2}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}=0,$ (17b)
$\sum_{k=1}^{K}\left[r_{k}+\frac{{\theta}_{k}^{2}h^{2}(t_{k})}{2\delta}\right]\leq P,{\forall t_{k}\in\mathcal{T}},{\forall k\in\mathcal{K}}.$ (17c)
In (17b), by replacing $r_{k}=\frac{{\theta}_{k}^{2}h^{2}(t_{k})}{2\delta(\beta-1)}$, we have
$\max_{\bm{t}}\quad$ $\sum_{k=1}^{K}{\left[\lambda g\left(\frac{\theta_{k}h^{2}(t_{k})}{\delta}\right)-\frac{{\theta}_{k}^{2}h^{2}(t_{k})}{2\delta(\beta-1)}-\frac{\theta_{k}^{2}h^{2}(t_{k})}{\delta}\right]},$ (18a)
s.t. $\sum_{k=1}^{K}\left[\frac{{\theta}_{k}^{2}h^{2}(t_{k})}{2\delta(\beta-1)}+\frac{\theta_{k}^{2}h^{2}(t_{k})}{\delta}\right]\leq P,{\forall t_{k}\in\mathcal{T}},{\forall k\in\mathcal{K}}.$ (18b)
Therefore, by utilizing exhaustive search, we can calculate optimal values of $h(t_{k})$ and $t_{k}$, then $r_{k}$, $R_{k}$, and $e_{k}$.
3.3 R3T Incentivization under Incomplete Information Case
We now introduce the contract game in an incomplete information setting. Clients choose contract items (i.e., effort, joining time, and corresponding reward) offered by the cloud server. While the cloud server aims to optimize its utility (e.g., global model performance), clients seek to maximize their effort-based incentives.
Lemma 1 (Monotonicity in CLPs).
The higher-type client should receive a higher reward $R$ when they join the training process early and contribute more effort. Otherwise, all clients would opt for higher-type contract items but join later and contribute less effort. Thus, under the IC constraints, if $\theta_{k}>\theta_{k^{\prime}},\forall k,k^{\prime}\in\mathcal{K}$, then the reward (including salary and bonus), effort, and joining time satisfy the following inequalities:
$R_{k}>R_{k^{\prime}},r_{k}\geq r_{k^{\prime}},B_{k}>B_{k^{\prime}},e_{k}>e_{k^{\prime}},t_{k}\leq t_{k^{\prime}}.$ (19)
Proof.
According to the IC constraints in (8), we have
$\frac{{\theta^{2}_{k}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}\geq\frac{{\theta^{2}_{k}}h^{2}(t_{k^{\prime}})}{2\delta}-(\beta-1)r_{k^{\prime}},$ (20)
$\frac{{\theta^{2}_{k^{\prime}}}h^{2}(t_{k^{\prime}})}{2\delta}-(\beta-1)r_{k^{\prime}}\geq\frac{{\theta^{2}_{k^{\prime}}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k},$ (21)
with $k,k^{\prime}\in\mathcal{K},k\neq k^{\prime}$. Summing up (20) and (21), we obtain
$\frac{h^{2}(t_{k})}{2\delta}({\theta^{2}_{k}}-{\theta^{2}}_{k^{\prime}})\geq\frac{h^{2}(t_{k^{\prime}})}{2\delta}({\theta^{2}_{k}}-{\theta^{2}}_{k^{\prime}}).$ (22)
As $\theta_{k}>\theta_{k^{\prime}}>0$, we have $({\theta_{k}}^{2}-{\theta}_{k^{\prime}}^{2})>0$. Dividing both sides by $({\theta_{k}}^{2}-{\theta}_{k^{\prime}}^{2})$, we have $h^{2}(t_{k})\geq h^{2}(t_{k^{\prime}})$ and as $h(t_{k})\geq 1,\forall{k\in\mathcal{K}}$, we conclude $h(t_{k})\geq h(t_{k^{\prime}})$. In addition, because $h(t_{k})$ is a strictly decreasing function of $t_{k}$, we can obtain $t_{k}\leq t_{k^{\prime}}$. Also, we prove if $h(t_{k})\geq h(t_{k^{\prime}})$, then $r_{k}\geq r_{k^{\prime}}$ under the IC constraints. Based on (21), we have
$(\beta-1)r_{k}-(\beta-1)r_{k^{\prime}}\geq\frac{{\theta^{2}_{k^{\prime}}}h^{2}(t_{k})}{2\delta}-\frac{{\theta^{2}_{k^{\prime}}}h^{2}(t_{k^{\prime}})}{2\delta},$
$(\beta-1)(r_{k}-r_{k^{\prime}})\geq\frac{{\theta}_{k^{\prime}}^{2}}{2\delta}[h^{2}(t_{k})-h^{2}(t_{k^{\prime}})]\geq 0.$ (23)
Since $\beta>1$ and $h(t_{k})\geq h(t_{k^{\prime}})$, thus $r_{k}\geq r_{k^{\prime}}$. Next, with the conditions of $\theta_{k}>\theta_{k^{\prime}},h(t_{k})\geq h(t_{k^{\prime}})$, and (11), we have $e_{k}>e_{k^{\prime}}$. As a result, we conclude $B_{k}>B_{k^{\prime}}$, thereby $R_{k}>R_{k^{\prime}}$ based on (1). This completes the proof. ∎
Lemma 2 (Client utility condition).
For any feasible contract item $(e_{k},t_{k},r_{k})$, the type-$k$ client’s utility satisfies
$\mathcal{U}_{n}(e_{1},t_{1},r_{1})<...<\mathcal{U}_{n}(e_{k},t_{k},r_{k})$ <...< …
$<\mathcal{U}_{n}(e_{K},t_{K},r_{K}).$ (24)
Proof.
According to Lemma 1, high-type clients who receive more rewards must contribute greater efforts and join as early as possible (i.e., $R_{k}>R_{k^{\prime}},e_{k}>e_{k^{\prime}}\text{ and }t_{k}\leq t_{k^{\prime}}$ are imposed together). Upon the conditions that ${\theta}_{k}>{\theta}_{k^{\prime}}$ and (11), we have
$\mathcal{U}_{n}(e_{k},t_{k},r_{k})=\frac{{\theta^{2}_{k}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}\geq\frac{{\theta^{2}_{k}}h^{2}(t_{k^{\prime}})}{2\delta}-(\beta-1)r_{k^{\prime}}$
$>\frac{{\theta^{2}_{k^{\prime}}}h^{2}(t_{k^{\prime}})}{2\delta}-(\beta-1)r_{k^{\prime}}=\mathcal{U}_{n}(e_{k^{\prime}},t_{k^{\prime}},r_{k^{\prime}}).$ (25)
This completes the proof. ∎
Lemma 3 (Simplify the IR constraints).
In the optimal contract, given that IC constraints in (8) are satisfied, the IR constraint for type-$1$ client is binding, i.e., $\frac{{\theta^{2}_{1}}h^{2}(t_{1})}{2\delta}-(\beta-1)r_{1}=0$.
Proof.
According to Lemma (2), we have
$\frac{{\theta^{2}_{k}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}$ $\geq\frac{{\theta^{2}_{1}}h^{2}(t_{1})}{2\delta}-(\beta-1)r_{1}\geq 0.$ (26)
As such, if the IR constraint of type-$1$ client is binding, that of all other types automatically holds. This completes the proof. ∎
The IR constraints in (7) can be replaced by Lemma (3), which indicates that this type-$1$ client gains or loses nothing in participating in the training process, while other client’s utilities are higher than that of the binding one. Next, we refer to [51] to introduce the following definitions for simplifying IC constraints:
Definition 5 (Conditions for IC constraints).
1) Downward incentive constraints (DICs) are IC constraints between client types $k$ and ${k^{\prime}}$, where $11\leq{k^{\prime}}\leq k-$ 1, 2) local downward incentive constraints (LDIC) is an IC constraint between client’s types $k$ and ${k^{\prime}},\text{where }{k^{\prime}}=k-1$, 3) upward incentive constraints (UICs) are IC constraints between client’s types $k$ and ${k^{\prime}},\text{where }{k}+1\leq k^{\prime}\leq{K}$, and 4) local upward incentive constraints (LUIC) is an IC constraint between client’s types $k$ and ${k^{\prime}},\text{where }k^{\prime}={k}+1$.
Lemma 4 (Simplify the IC constraints).
In the optimal contract, we can replace the IC constraints in (8) by
$\frac{{\theta^{2}_{k}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}=\frac{{\theta^{2}_{k}}h^{2}(t_{k-1})}{2\delta}-(\beta-1)r_{k-1}.$ (27)
Proof.
We consider three client types where ${\theta}_{k-1}<{\theta}_{k}<{\theta}_{k+1},\forall k\in\mathcal{K}$. Under IC constraints, we have two LDICs inequalities as follows
$\frac{{\theta^{2}_{k+1}}h^{2}(t_{k+1})}{2\delta}-(\beta-1)r_{k+1}\geq\frac{{\theta^{2}_{k+1}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k},$ (28)
$\frac{{\theta^{2}_{k}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}\geq\frac{{\theta^{2}_{k}}h^{2}(t_{k-1})}{2\delta}-(\beta-1)r_{k-1}.$ (29)
By virtue of monotonicity in Lemma 1, the condition of $\theta$, and (29), we have
$\frac{{\theta}_{k+1}^{2}}{2\delta}\left[h^{2}(t_{k})-h^{2}(t_{k-1})\right]$ $\geq\frac{{\theta}_{k}^{2}}{2\delta}\left[h^{2}(t_{k})-h^{2}(t_{k-1})\right]$
$\geq(\beta-1)(r_{k}-r_{k-1}).$ (30)
As a result, we convert (28) as
$\frac{{\theta^{2}_{k+1}}h^{2}(t_{k+1})}{2\delta}-(\beta-1)r_{k+1}$ $\geq\frac{{\theta^{2}_{k+1}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}$
$\geq\frac{{\theta}_{k+1}^{2}h^{2}(t_{k-1})}{2\delta}-(\beta-1)r_{k-1}.$ (31)
Therefore, we conclude that LDIC has transitivity, i.e., if there is an LDIC between type-$(k-1)$ and type-$k$ clients, then we can further extend incentive constraints from type-$(k-1)$ client to type-$1$ client, that is, all DICs hold, i.e.,
$\frac{{\theta^{2}_{k+1}}h^{2}(t_{k+1})}{2\delta}-(\beta-1)r_{k+1}$ $\geq\frac{{\theta}_{k+1}^{2}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}$
$\geq\frac{{\theta}_{k+1}^{2}h^{2}(t_{k-1})}{2\delta}-(\beta-1)r_{k-1}$
$\geq$...≥ …
$\geq\frac{{\theta}_{k+1}^{2}h^{2}(t_{1})}{2\delta}-(\beta-1)r_{1}.$ (32)
Thus, we have proved that with the LDIC, all DICs hold. Similarly, with the LUIC, all UICs are satisfied. We can write a generalization of DICs and UICs, respectively, as follows
$\frac{{\theta^{2}_{k}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}$ $\geq\frac{{\theta}_{k}^{2}h^{2}(t_{k{{}^{\prime}}})}{2\delta}-(\beta-1)r_{k{{}^{\prime}}},$
$1\leq k{{}^{\prime}}<k\leq K,$
$\frac{{\theta^{2}_{k}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}$ $\geq\frac{{\theta}_{k}^{2}h^{2}(t_{k{{}^{\prime}}})}{2\delta}-(\beta-1)r_{k{{}^{\prime}}},$
$1\leq k<k{{}^{\prime}}\leq K.$ (33)
Therefore, the IC constraints are reduced to
$\frac{{\theta^{2}_{k}}h^{2}(t_{k})}{2\delta}-(\beta-1)r_{k}$ $\geq\frac{{\theta}_{k}^{2}h^{2}(t_{k-1})}{2\delta}-(\beta-1)r_{k-1},$ (34)
which can replace IC constraints in (8). This completes the proof. ∎
Input: A set of clients $\mathcal{N}$, number of training round $T$, learning rate $\eta$, local loss function $F$, initial global parameter $w^{0}$, bonus unit coefficient $\vartheta$.
Output: Final global model parameter $w^{T}$.
Data: Training, validation, and test set $\mathcal{D}$.
1
2for $t$ = 1, 2,…, T do
[…]
4      Obtain an optimal contract set $\bm{\phi^{*}}$ via (18) or (38).
5      Send the set $\bm{\phi^{*}}=\{\bm{e^{*}},\bm{t^{*}},\bm{r^{*}}\}=\{\phi_{k}^{*}\}_{k\in\mathcal{K}}$ to $\mathcal{N}$.
6      Select a client subset $S_{n}^{t}$ from $\mathcal{N}$.
[…]
8      for each client $n\in S_{n}^{t}$ in parallel do
9             Sign contract items $\{\phi_{k}^{*}\}_{k\in\mathcal{K}}$.
10            if $t^{*}_{k}=t$ then
11                   Download $w^{t-1}$.
12                  Adjust $\mathcal{D}_{k}$ according to $e^{*}_{k}$.
13                  $w^{t}_{k}\leftarrow w^{t-1}_{k}-\eta\triangledown{F}_{k}(w^{t-1}_{k},\mathcal{D}_{k})$
14
[…]
16      Cross-validate and record {$ID_{k}$, $hash(w^{t}_{k})$, $\phi_{k}^{*}$}.
[…]
18      if $t\in{\text{CLPs}}$ then
19
20            $|S_{n}^{t+1}|\leftarrow\text{min}\{2|S_{n}^{t}|,N\}$
21            $h(t^{*}_{k})=h(t)=1+\frac{\vartheta}{\ln(2t)}$
22      else
23             $|S_{n}^{t+1}|\leftarrow\text{max}\{\frac{1}{2}|S_{n}^{t}|,\frac{1}{2}N\}$
24            $h(t^{*}_{k})=1$
25
26      $w^{t}\leftarrow\sum_{k\in\mathcal{K}}{\frac{e^{*}_{k}}{\cup_{k\in\mathcal{K}}e^{*}_{k}}}w^{t}_{k}$
[…]
28      Record {$h(t^{*}_{k}),w^{t}$}.
[…]
30Calculate and allocate $R_{k}^{*}$ to type-$k$ clients via (1).
Return $w^{T}$.
Algorithm 1 Game and Training Process of R3T
Theorem 2.
For any type-$k$ client, the contract salary satisfies
$r_{k}=\sum_{i=1}^{k}\frac{\theta^{2}_{i}[h^{2}(t_{i})-h^{2}(t_{i-1})]}{2\delta(\beta-1)}.$
Proof.
Based on Lemma 3, the salary of type-$1$ client can be calculated as follows
$r_{1}=\frac{\theta^{2}_{1}h^{2}(t_{1})}{2\delta(\beta-1)}.$ (35)
Then, according to Lemma 4 and (35), we can further derive
$r_{2}=r_{1}+\frac{\theta^{2}_{2}[h^{2}(t_{2})-h^{2}(t_{1})]}{2\delta(\beta-1)}.$ (36)
By iterating these steps, we can obtain
$r_{k}$ $=r_{k-1}+\frac{\theta^{2}_{k}[h^{2}(t_{k})-h^{2}(t_{k-1})]}{2\delta(\beta-1)}$
$=\frac{\theta^{2}_{k}[h^{2}(t_{k})-h^{2}(t_{k-1})]}{2\delta(\beta-1)}+...+\frac{\theta^{2}_{1}h^{2}(t_{1})}{2\delta(\beta-1)}$
$=\sum_{i=1}^{k}\frac{\theta^{2}_{i}[h^{2}(t_{i})-h^{2}(t_{i-1})]}{2\delta(\beta-1)}.$ (37)
Note that at $i=1$, we have $h(t_{0})=0$ because $t_{0}$ is the time before the training process starts. This completes the proof. ∎
By using the simplified constraints above and substituting $e_{k}$ and $r_{k}$ with $h(t_{k})$, the final form of the optimization problem is as follows
$\max_{\bm{t}}\quad$ $\sum_{k=1}^{K}\Big{[}\lambda g(\frac{\theta_{k}h^{2}(t_{k})}{\delta})-\sum_{i=1}^{k}\frac{\theta^{2}_{i}[h^{2}(t_{i})-h^{2}(t_{i-1})]}{2\delta(\beta-1)}$
[…]
s.t. $\sum_{k=1}^{K}\left[\sum_{i=1}^{k}\frac{\theta^{2}_{i}[h^{2}(t_{i})-h^{2}(t_{i-1})]}{2\delta(\beta-1)}+\frac{\theta_{k}^{2}h^{2}(t_{k})}{\delta}\right]\leq P,$ (38b)
(19)$,{\forall t_{k}\in\mathcal{T}},{\forall k\in\mathcal{K}}.$ (38c)
Using an exhaustive search algorithm, we obtain the optimal $t_{k}$ and $h(t_{k})$, then $r_{k}$ and $R_{k}$, and finally $e_{k}$ without considering the monotonicity condition in (19). Then, we check if the solution satisfies the monotonicity condition in (19).
Based on the analyses and proofs above, we now analyze monotonicity conditions in the non-CLPs.
Lemma 5 (Monotonicity in non-CLPs).
In the non-CLPs, R3T still ensures incentive fairness by enabling a higher-type client to receive a higher reward. Thus, to satisfy the IC constraints in (8), if $\theta_{k}>\theta_{k^{\prime}},\forall k,k^{\prime}\in\mathcal{K}$, then the reward, salary, and effort must adhere to the following inequalities:
$R_{k}>R_{k^{\prime}},r_{k}=r_{k^{\prime}},e_{k}>e_{k^{\prime}}.$ (39)
Proof.
According to (1), $h(t_{k})=1$ and the impact of efforts in non-CLPs is not of importance as CLPs, thereby $h(t_{k^{\prime}})=h(t_{k})=1,\forall{t_{k},t_{k^{\prime}}\notin\text{CLPs}}$ This leads to $e_{k}>e_{k^{\prime}}$ based on (11). Moreover, via (3.3) and $h(t_{k^{\prime}})=h(t_{k})$, we can conclude that $r_{k}=r_{k^{\prime}}=r_{1}=\frac{\theta^{2}_{1}h^{2}(t_{1})}{2\delta(\beta-1)}$. Next, with the proved conditions (i.e., $h(t_{k})=h(t_{k^{\prime}})\text{ and }r_{k}=r_{k^{\prime}}$), and the expression (1), we have $R_{k}>R_{k^{\prime}}$. This completes the proof. ∎
Remark 1.
In non-CLPs, the impact of the client’s contribution is less significant compared to CLPs. As Lemma 5 outlines, although all clients receive a fixed salary regardless of their types and do not receive weighted bonuses (i.e., $h(t_{k})=1$) for specific joining times, R3T ensures the fairness of their total rewards relative to their corresponding efforts.
Remark 2.
R3T handles the information asymmetry between clients and the cloud server via a self-revelation principle by designing a menu of contract items to incentivize clients to disclose their private properties and an incentive compatibility constraint to encourage clients to truthfully choose and report their type-aligned contract items to get the best utilities. R3T guarantees the fairness in incentive distribution in both CLPs and non-CLPs as shown in Lemma 1 and Lemma 5.
The R3T’s game and training processes in $T$ training rounds are depicted in Algorithm 1. We have a tuple of $(e^{*}_{k},t^{*}_{k},r^{*}_{k})$ being the optimal strategy of the type-$k$ client (line 1). CLPs can be identified by using federated gradient norm metric (line 1) [11]. To address the issue of insufficient client participation and its detrimental effect on final model performance, R3T increases the number of clients during the early training phase (line 1) by offering an attractive contract set. Besides, to reduce communication overhead and budget exhaustion, R3T decreases the number of clients in non-CLPs (line 1).
4 Experimental Results
(a) Client’s effort.
(b) Client’s fixed salary.
(c) Client’s bonus.
(d) Client’s reward.
(e) Client utility.
(f) Cloud utility.
Figure 4: The impact of unit effort cost $\delta$ on the client’s effort, fixed salary, bonus, total reward, client utility, and cloud utility.
In this section, we conduct extensive experiments to analyze the feasibility and efficiency of R3T. Then, we evaluate R3T’s proof of concept on prominent benchmark datasets.
TABLE I: Experimental Parameters
Param Value Param Value Param Value
$\lambda_{\text{{CLPs}}}$ 21 $\lambda_{\text{{non-CLPs}}}$ 20 $\theta_{k}$ U(1, 2)
$T$ 25 N 15 K 10
$C$ 2.4 $\beta$ 3 $t_{\text{{non-CLPs}}}$ [11, $T$]
4.1 Experiment Setup
Simulation parameter settings. Without loss of generality, our experimental settings refer to the setups in [19, 52]. The key parameters used in the simulations are shown in Table I.
Proof of concept. We develop a system to implement a proof of concept of R3T, leveraging smart contracts on an Ethereum blockchain using Ganache [53]. The Flower framework [54] is integrated with the blockchain to facilitate federated learning experiments for model training and evaluation. The system is deployed on an Ubuntu 22.04 system equipped with an Intel Core i7-1365U CPU@1.80 GHz and 32GB of memory. To enable interaction between clients, the cloud server, and the blockchain smart contracts, we utilize the Web3 API [55] and FastAPI frameworks [56]. Also, we develop two smart contracts for federated training and storage, and reward calculation using Solidity programming language.
Model and Datasets: We use a deep neural network with four layers comprising three convolutional neural layers with dropout to prevent overfitting, batch normalization, and max pooling for reducing spatial dimensions, and one dense layer. For model training and testing, CIFAR-10 [57] and Fashion-MNIST [58] datasets are used in a non-identical and independent (non-IID) setting. We consider a highly heterogeneous setting in which the number of data samples and data labels among clients are unbalanced by sampling $p\sim\textit{Dir}{(\alpha)}$, where $\alpha=\{0.1,0.4,0.7\}$ is the parameter of Dirichlet distribution. Besides, the learning rate $\eta=$ is 3e-4 and the batch size is 32.
Benchmarks. We consider the following benchmarks in CLPs and non-CLPs for comparison with our proposed methods, which are labeled as R3T in Complete Information Case (R3T-CIC) and R3T in Incomplete Information Case (R3T-IIC).
  • Contract theory Without considering Time under CIC (CTWT-CIC). Time property is not considered in the existing work (e.g., [51, 31, 19, 23]), that is, $t$ is excluded from the contract item. All training time slots are treated equally.
  • Contract theory Without considering Time under IIC (CTWT-IIC). Similar to CTWT-CIC, this method does not consider temporal importance. In addition, the cloud server only knows the distribution of clients’ types.
  • Linear Pricing. Under IIC, the cloud server determines a unit price $C$ for the client’s effort $e$ [21].
  • Conventional Federated Learning (CFL). FL methods integrating conventional incentive mechanisms (e.g., [23, 31, 32]), which do not consider CLPs, are used to compare with R3T proof of concept.
(a) Under CIC.
(b) Under IIC.
Figure 5: The impact of cloud budget $P$ on cloud utility per contract item and total cloud utility.
4.2 Contract Feasibility
We first examine the impacts of varying unit effort cost $\delta$ and cloud budget $P$ on R3T over $N=10$ clients in R3T during the initial training round of CLPs. Then, we evaluate cloud utility, effort, reward, and client utility against contract items (i.e., client types) in R3T during the initial training round of CLPs and non-CLPs starting at $t=1$ and $t=11$, correspondingly, where we set the cloud budget $P=60$ per round and $\delta=1$.
Impacts of different unit effort costs. In Figs.  4(4(a))-(4(f)), under CIC and IIC in CLPs [^9] We focus on one global training round in CLPs here, but analogous trends observed in non-CLPs., the client’s fixed salary, bonus, effort, total reward, utility, and cloud utility have shown decreasing trends per contract item along with the increase of unit effort cost $\delta$, where we set $\delta=\{0.2,0.6,1\}$. In other words, clients are not willing to join and contribute efforts when $\delta$ becomes large. For example, given the tenth contract item under the IIC cases, the values of the client’s effort units are around {9.60, 3.20, 1.92} corresponding to $\delta=\{0.2,0.6,1\}$. Besides, under CIC, where the cloud server possesses complete client information, the cloud can fine-tune and optimize its contract items to incentivize clients to join early and contribute more effort. This results in higher cloud utility while offering clients higher fixed salaries and bonuses, compared to IIC (e.g., 232.28 vs. 211.13, 5.96 vs. 1.01, and 119.21 vs. 18.47 given the tenth contract item with $\delta=0.2$, respectively). In addition, due to being aware of client information, the cloud server exploits the client’s efforts as much as possible while guaranteeing IR constraints (c.f. Theorem 1) in CIC, leading client utilities over contract items equal to zero (c.f., Fig. 4(4(e))).
(a) Cloud utility.
(b) Effort.
(c) Reward.
Figure 6: Performance evaluation of R3T per training round.
Impacts of different cloud budgets. Figs.  5(5(a))-(5(b)) show the effect of varying cloud budgets $P$ on the cloud utility contributed by type-$k$ clients (i.e., $k$-th contract items) in a global training round under both CIC and IIC in CLPs. As $P$ increases, more clients can be incentivized to participate in and allocate efforts, resulting in greater total cloud utility over all joined clients. For instance, at $P=60$, the total cloud utility of all type-$k$ clients reaches approximately 375 in CIC and 312 in IIC, compared to 266 in CIC and 311 in IIC at $P=30$, respectively. Especially at low budgets (i.e., $P=\{10,20,30\}$), the total cloud utility under IIC exceeds that under CIC. This is supported by Assumption 1 and the fact that the cloud server is aware of clients’ types in CIC and only the distribution of clients’ types in IIC, thereby different optimal contract designs. Consequently, in CIC, the cloud server may adopt a greedy allocation strategy to exhaust efforts from higher-type clients (c.f. Theorem 1), but due to a limited budget, it leads to fewer participating clients and thus lower total cloud utility compared with that of IIC. Conversely, when the budget is huge enough (i.e., $P\geq 40$), the total cloud utility in CIC is far larger than that in IIC (i.e., 375 versus 311 at $P=60$).
Figure 7: Cloud utility with different number of contract items.
Performance evaluation. We consider cloud utility, client’s effort, and client’s reward according to contract items in a training round in CLPs and non-CLPs under conditions of CIC and IIC as shown in Figs. 6(6(a))-(6(c)). In Fig. 6(6(a)), the cloud server achieves the maximum utilities in both CLPs and non-CLPs in CIC compared to IIC. This outcome occurs because the cloud server has complete knowledge of the client’s capabilities, enabling it to fully leverage clients’ effort and time without breaching IR constraints, thereby maximizing utility as per Theorem 1. Conversely, under IIC, the cloud server does not obtain complete information but only distribution of the client’s type; therefore the cloud server’s utility remains constrained by the utility attainable under CIC. Figs. 6(6(b))-(6(c)) contrast efforts and rewards (comprising of salaries and bonuses) associated with different contract items. These results validate the monotonicity conditions, which are detailed in Lemma 1 and Lemma 5, and Remark 1. As can be seen from the figures, higher-type clients contribute more effort and, consequently, receive higher rewards.
Scalability. In Fig. 7, we study the influence of the number of contract items among clients on the cloud utility. As the number of contract items increases, the cloud server can enhance its utility by enabling and motivating more higher-type clients to participate early and contribute more effort. R3T shows its scalability in diverse and expansive training environments (e.g., Internet of Things). However, expanding the number of contract items also presents a challenge of budget constraints for the cloud server.
Figure 8: Client utility under different contract items.
Contract feasibility. In Fig. 8, the utilities of four client types are compared when they select the same contract item. Their utilities follow the inequality $U_{1}<U_{4}<U_{7}<U_{10}$ and each client can achieve maximum utility if and only if they select the contract item designed for their types, corroborating the results presented in Lemma 2 and Lemma 1, respectively, which explain the IC constraint. This contract design ensures that clients’ types are automatically disclosed to the cloud server upon contract selection, effectively addressing information asymmetry between clients and the cloud server. Furthermore, clients receive non-negative utilities when they choose the corresponding contract items with their types, which validates the IR constraint.
(a) Cloud utility.
(b) Total cloud utility.
(c) Economic efficiency.
Figure 9: Cloud gain over $T$ training rounds.
(a) Client utility.
(b) Client’s reward.
Figure 10: Client gain over $T$ training rounds.
4.3 Contract Efficiency
Fig. 9 and Fig. 10 examine the efficiency of R3T with CTWT and linear pricing with the same cloud budget $P=60$, where the number of clients is $N=15$ and unit effort cost $\delta=1$. In this comparison, we assume that CLPs span from $t=1$ to $t=10$, and non-CLPs begin from $t=11$ to the end.
Cloud utility. Fig. 9(9(a)) shows cloud utilities in the CLPs and non-CLPs. R3T demonstrably outperforms other methods by effectively attracting the highest-quality contributions in the CLPs. Consequently, the cloud server maintains a significantly higher utility level compared with other approaches during the CLPs, which determine the final model performance. Although the cloud utility of R3T shows a gradual decrease, which becomes more pronounced once the learning and training process transitions to the non-CLPs (i.e., $t\geq 11$), the total R3T’s cloud utility still achieves higher results in comparison with conventional incentive benchmarks, as shown in Fig. 9(9(b)). For example, R3T showcases cloud utility up to approximately 7.2% and 23.1% higher in the IIC compared to CTWT-IIC and linear pricing methods, respectively.
Moreover, R3T demonstrates higher economic efficiency compared to conventional benchmarks, as shown in Fig. 9(9(c)). Economic efficiency reflects the effectiveness of the cloud server in calculating and distributing the total amount of reward to clients based on their efforts and joining times. For instance, R3T spends approximately 20% and 96% less on the amount of reward compared with CTWT and linear pricing, respectively, while still reaching higher final cloud utility in the IIC setting. This highlights the importance of providing the right reward at the right time.
Client gain. As the importance of training rounds diminishes in non-CLPs, client utilities trend downward in R3T, as shown in Fig. 10(10(a)). Meanwhile, the CTWT methods, which treat training rounds equally and randomly select a number of clients per round without early participation incentives, maintain stable client utilities - roughly 7.5 in CTWT-IIC methods and 0 in CTWT-CIC methods as indicated in Theorem 1 throughout the $T$ training rounds. The linear pricing method exhibits the lowest cloud utility but the highest client utility due to its fixed price per unit of client effort. This fixed pricing approach limits information gathering for the cloud and prevents optimal price and reward adjustments to match varying client capabilities, efforts, and costs.
Fig. 10(10(b)) shows the total client’s rewards offered in $T$ training rounds. While R3T allocates a significant amount of rewards to attract high-quality clients during the CLPs, the amount of rewards of other benchmarks remains generally stable in every training round. This again stems from the strategic allocation of rewards, prioritizing efforts made earlier in the training process in R3T, and further explaining the economic efficiency of the cloud server.
4.4 Proof of Concept of R3T
In Fig. 11, we introduce a proof of concept of R3T. Clients with higher IDs are assigned higher training capabilities, characterized by larger data sizes and increased data diversity. We assume all clients have identical system capabilities and act rationally and strategically under Assumption 1. Besides, CLPs end at training round 31. Clients register on the blockchain and their information (e.g., IDs, addresses, efforts, and joined training rounds) is recorded throughout the training. Therefore, non-repudiation and transparency are guaranteed thanks to blockchain properties.
Figure 11: R3T’s proof of concept. The system includes the Ethereum blockchain, online IPFS storage to address scalability issues, a cloud dashboard and ten clients’ dashboards.
(a) CIFAR-10.
(b) Fashion-MNIST.
Figure 12: Test accuracy of R3T and CFL.
Convergence rate. As illustrated in Fig. 12, R3T demonstrates its efficiency by surpassing CFL in reducing convergence rate under non-IID settings for both CIFAR-10 and Fashion-MNIST datasets. Particularly, R3T achieves a significant increase during the initial FL training phase. This improvement is attributed to its strategy of attracting the highest-quality clients by offering an optimal set of contract items, which includes timely and substantial rewards, early in the training process. For example, R3T tends to reach the final target accuracy in just 20 rounds, compared with approximately 61 rounds for CFL on the datasets.
Long-term performance. For different values of the parameter of Dirichlet distribution $\alpha=\{0.4,0.7\}$ in both datasets, both R3T and CFL maintain stable test accuracies after convergence, even though R3T selects a smaller number of participating clients in these periods (c.f., in line 1 of Algorithm 1). Under a highly non-IID setting (i.e., $\alpha=0.1$), R3T demonstrates more sustainable performance compared to CFL. In CFL, the practice of treating clients’ contributions and efforts equally across all training rounds, especially under a significant quantity and label distribution skew, detrimentally affects the long-term performance of the FL model, regardless of subsequent efforts. Besides, although R3T shows a slight drop in accuracy after CLPs due to the reduced number of selected clients, the overall test accuracy remains stable around the highest bound. These observations highlight the significant impact of attracting high-quality clients early, saving around 200%-300% in the total training time and 5.2%-47.6% in the total number of clients to achieve a close targeted learning performance.
5 Conclusion
In this work, we studied a time-aware incentive mechanism implemented on blockchain smart contracts for FL, referred to as R3T. By characterizing the interactions between rational client’s time and system capabilities, efforts, joining time, reward, and their impact on cloud utility, we proposed a contract design that incorporates the temporal factor and then derived the optimal incentive strategies for the cloud. R3T transparently provides the right reward at the right time, thereby attracting the highest-quality contributions in the early training phase, which is crucial in determining the final model performance under information symmetry and information asymmetry conditions. Simulation and proof-of-concept experimental results demonstrate the superiority of R3T over incentive mechanism benchmarks in improving cloud utility and economic efficiency, reducing convergence time and the total number of participating clients, while achieving comparable model performance.
In future work, we will consider impacts of free-rider issues and malicious attacks on the efficiency of R3T, especially in CLPs. Besides, it would be interesting to evaluate the efficiency and scalability of R3T on different blockchain platforms integrating FL systems, where the entire system is operated without relying on the cloud server.